# A Survey on Knowledge Distillation of Large Language Models

### 20 Feb 2024 v1

@@ 4.4.6 Code
LLMs, trained on extensive corpora containing code, are highlighted for their proficiency in code-related tasks. Their capabilities extend beyond direct code generation to include the provision of external knowledge and data, which is crucial in distilling their expertise into smaller, more efficient models. Several works have successfully distilled code knowledge from LLMs into those compact and specialized code models Chaudhary (2023); Rozière et al. (2023); Gunasekar et al. (2023); Wei et al. (2023); Chen et al. (2023a); Liu et al. (2023e); Yu et al. (2024); Jain et al. (2023); Su and McMillan (2023); Guo et al. (2023d).
@@
However, these models primarily target the code generation task, lacking generalizability across a broader range of code-related tasks. To address this issue, MFTCoder Liu et al. (2023e) utilizes self-instruct to distill diverse code data from teacher LLMs for various tasks, such as code completion and text-to-code generation, training a student model via multi-task learning. WaveCoder Yu et al. (2024), in contrast, creates a comprehensive instruction tuning dataset covering four universal code-related tasks distilled from GPT-3.5-turbo. WaveCoder first selects a diverse coreset of raw data using the KCenterGreedy Sener and Savarese (2018) clustering method, then employs the teacher LLM for generating task definitions and outputs. The teacher model also plays a role in evaluating and filtering this data. Notably, WaveCoder demonstrates superior generalization across different code-related tasks compared to other open-source models.
@@ 4.5 Multi-Modality
Multimodal Large Language Models (MLLMs) surpass traditional language-only LLMs by understanding and processing information across multiple modalities, more closely mirroring human perception and enabling a broader range of real-world applications. There is a growing trend towards developing MLLMs that follow multimodal instructions, facilitating tasks with enhanced levels of interactivity. To address the scarcity of multimodal instruction-following data and to harness the commonsense and world knowledge embedded in teacher LLMs, numerous studies have focused on multimodal knowledge distillation from LLMs Liu et al. (2023b); Zhao et al. (2023b); Wang et al. (2023e); Chen et al. (2023c); Park et al. (2023); Pi et al. (2023); Zhao et al. (2023c); Liu et al. (2023f); Wu et al. (2023b); Luo et al. (2023d); Jiang et al. (2023d); Li et al. (2023c); Xu et al. (2023e).
Vision-Language. In the vision-language domain, LLaVA Liu et al. (2023b) pioneers the extension of the Self-Instruct approach from the language to the multimodal field. It translates images into textual descriptions, including captions and bounding boxes, and distills GPT-4 for generating new data in the context of seed examples. This approach creates a LLaVA-Instruct-150k dataset, which serves as the foundation for further developments like LLaVA-1.5 Liu et al. (2023l) and GPT4ROI Zhang et al. (2023e), enhancing the instruction-following capabilities of MLLMs. To expand the dataset’s scale, SVIT Zhao et al. (2023b) introduces a 4.2 million image dataset, distilled from GPT-4 by leveraging manual image annotations. It employs a novel data recipe to select an informative, diverse, and balanced subset of training data. LVIS-Instruct4V Wang et al. (2023e) leverages GPT-4V OpenAI (2023), a powerful large multimodal model, as a teacher to distill a more accurate and context-aware instruction-following dataset, focusing on fine-grained understanding. Further advancements include integrating specific region referencing in image-based instruction following. For instance, Shikra Chen et al. (2023c) uses GPT-4 to distill referential question-answer pairs from the Flickr30K Plummer et al. (2015) dataset, enhancing the understanding of referential regions within images. LSKD Park et al. (2023) introduces localized references to specific image regions, prompting the teacher LLM to generate commonsense inferences about these areas. To enhance the visual instruction tuning pipeline with text-rich images, LLaVAR Zhang et al. (2023d) employs the text-only GPT-4 as a teacher, using recognized texts and image captions to generate 16K conversation pairs for text-rich images. The resultant student MLLM demonstrates enhanced interaction skills in content that combines both text and imagery.
@@
MedAlpaca Han et al. (2023), PMC-LLaMA Wu et al. (2023e), DISC-MedLLM Bao et al. (2023a), leaf, text width=33em ] ] [ Finance [ XuanYuan Zhang and Yang (2023), leaf, text width=33em ] ] [ Science [ DARWIN Xie et al. (2023a), SciGLM Zhang et al. (2024), WizardMath Luo et al. (2023c),
@@ 5.4 Science
Mathematics. The application of LLMs within the sub-domain of mathematics heralds a transformative era in mathematical research, education, and problem-solving (Azerbayev et al., 2023; Yu et al., 2023b). LLMs in mathematics facilitate the exploration and understanding of complex mathematical theories and problems by providing intuitive explanations, proofs, and solutions that can bridge the gap between advanced mathematical concepts and learners at various levels. These models have shown potential in conjecturing new mathematical theorems and patterns, thus opening new avenues for research and discovery that might not have been readily accessible to humans alone. In education, they serve as personalized tutors, offering students step-by-step guidance through mathematical problems and adapting explanations to the learner’s level of understanding. This democratizes access to high-quality mathematical education and fosters a deeper appreciation and understanding of mathematics among a broader audience. By enhancing collaborative efforts through the generation of new ideas and the simplification of complex concepts, LLMs are poised to significantly advance the field of mathematics, making it more accessible, efficient, and innovative. WizardMath (Luo et al., 2023c) enhances the mathematical reasoning capabilities of Llama-2 by applying the novel Reinforcement Learning from Evol-Instruct Feedback (RLEIF) method, significantly outperforming other open-source LLMs on the GSM8k and MATH benchmarks, as well as surpassing several closed-source LLMs including ChatGPT-3.5 and Minerva. MAmmoTH (Yue et al., 2023a) is a series of open-source LLMS specifically developed for general math problem-solving, achieving superior performance on nine mathematical reasoning datasets. Utilizing a novel instruction tuning dataset called MathInstruct, which combines chain-of-thought and program-of-thought rationales, MAmmoTH models demonstrate substantial improvements over existing models. TORA (Gou et al., 2024), a series of Tool-integrated Reasoning Agents, significantly advances mathematical problem-solving by combining natural language reasoning with the use of external computational tools. It markedly outperforms existing open-source models on 10 mathematical reasoning datasets, showcasing notable improvements over both rationale-based and program-based approaches, and introduces innovative training techniques such as output space shaping to enhance model reasoning capabilities. G-LLaVA (Gao et al., 2023c) introduces a significant advancement in geometric problem-solving for LLMs by leveraging a multimodal approach that combines text and image data. This model, utilizing the Geo170K dataset comprising over 170,000 geometric image-caption and question-answer pairs, demonstrates remarkable improvements over GPT-4V on the MathVista benchmark.

### 23 Feb 2024 v2

@@ 4.4.6 Code
LLMs, trained on extensive corpora containing code, are highlighted for their proficiency in code-related tasks. Their capabilities extend beyond direct code generation to include the provision of external knowledge and data, which is crucial in distilling their expertise into smaller, more efficient models. Several works have successfully distilled code knowledge from LLMs into those compact and specialized code models Chaudhary (2023); Rozière et al. (2023); Gunasekar et al. (2023); Wei et al. (2023); Chen et al. (2023a); Liu et al. (2023d); Yu et al. (2024); Jain et al. (2023); Su and McMillan (2023); Guo et al. (2023d).
@@
However, these models primarily target the code generation task, lacking generalizability across a broader range of code-related tasks. To address this issue, MFTCoder Liu et al. (2023d) utilizes self-instruct to distill diverse code data from teacher LLMs for various tasks, such as code completion and text-to-code generation, training a student model via multi-task learning. WaveCoder Yu et al. (2024), in contrast, creates a comprehensive instruction tuning dataset covering four universal code-related tasks distilled from GPT-3.5-turbo. WaveCoder first selects a diverse coreset of raw data using the KCenterGreedy Sener and Savarese (2018) clustering method, then employs the teacher LLM for generating task definitions and outputs. The teacher model also plays a role in evaluating and filtering this data. Notably, WaveCoder demonstrates superior generalization across different code-related tasks compared to other open-source models.
@@ 4.5 Multi-Modality
Multimodal Large Language Models (MLLMs) surpass traditional language-only LLMs by understanding and processing information across multiple modalities, more closely mirroring human perception and enabling a broader range of real-world applications. There is a growing trend towards developing MLLMs that follow multimodal instructions, facilitating tasks with enhanced levels of interactivity. To address the scarcity of multimodal instruction-following data and to harness the commonsense and world knowledge embedded in teacher LLMs, numerous studies have focused on multimodal knowledge distillation from LLMs Liu et al. (2023e); Zhao et al. (2023b); Wang et al. (2023e); Chen et al. (2023c); Park et al. (2023); Pi et al. (2023); Zhao et al. (2023c); Liu et al. (2023f); Wu et al. (2023b); Luo et al. (2023d); Jiang et al. (2023d); Li et al. (2023c); Xu et al. (2023e).
Vision-Language. In the vision-language domain, LLaVA Liu et al. (2023e) pioneers the extension of the Self-Instruct approach from the language to the multimodal field. It translates images into textual descriptions, including captions and bounding boxes, and distills GPT-4 for generating new data in the context of seed examples. This approach creates a LLaVA-Instruct-150k dataset, which serves as the foundation for further developments like LLaVA-1.5 Liu et al. (2023l) and GPT4ROI Zhang et al. (2023e), enhancing the instruction-following capabilities of MLLMs. To expand the dataset’s scale, SVIT Zhao et al. (2023b) introduces a 4.2 million image dataset, distilled from GPT-4 by leveraging manual image annotations. It employs a novel data recipe to select an informative, diverse, and balanced subset of training data. LVIS-Instruct4V Wang et al. (2023e) leverages GPT-4V OpenAI (2023), a powerful large multimodal model, as a teacher to distill a more accurate and context-aware instruction-following dataset, focusing on fine-grained understanding. Further advancements include integrating specific region referencing in image-based instruction following. For instance, Shikra Chen et al. (2023c) uses GPT-4 to distill referential question-answer pairs from the Flickr30K Plummer et al. (2015) dataset, enhancing the understanding of referential regions within images. LSKD Park et al. (2023) introduces localized references to specific image regions, prompting the teacher LLM to generate commonsense inferences about these areas. To enhance the visual instruction tuning pipeline with text-rich images, LLaVAR Zhang et al. (2023d) employs the text-only GPT-4 as a teacher, using recognized texts and image captions to generate 16K conversation pairs for text-rich images. The resultant student MLLM demonstrates enhanced interaction skills in content that combines both text and imagery.
@@
MedAlpaca Han et al. (2023), PMC-LLaMA Wu et al. (2023e), DISC-MedLLM Bao et al. (2023a), leaf, text width=33em ] ] [ Finance [ XuanYuan Zhang and Yang (2023), leaf, text width=33em ] ] [ Science [ DARWIN Xie et al. (2023a), SciGLM Zhang et al. (2024), WizardMath Luo et al. (2023b),
@@ 5.4 Science
Mathematics. The application of LLMs within the sub-domain of mathematics heralds a transformative era in mathematical research, education, and problem-solving (Azerbayev et al., 2023; Yu et al., 2023b). LLMs in mathematics facilitate the exploration and understanding of complex mathematical theories and problems by providing intuitive explanations, proofs, and solutions that can bridge the gap between advanced mathematical concepts and learners at various levels. These models have shown potential in conjecturing new mathematical theorems and patterns, thus opening new avenues for research and discovery that might not have been readily accessible to humans alone. In education, they serve as personalized tutors, offering students step-by-step guidance through mathematical problems and adapting explanations to the learner’s level of understanding. This democratizes access to high-quality mathematical education and fosters a deeper appreciation and understanding of mathematics among a broader audience. By enhancing collaborative efforts through the generation of new ideas and the simplification of complex concepts, LLMs are poised to significantly advance the field of mathematics, making it more accessible, efficient, and innovative. WizardMath (Luo et al., 2023b) enhances the mathematical reasoning capabilities of Llama-2 by applying the novel Reinforcement Learning from Evol-Instruct Feedback (RLEIF) method, significantly outperforming other open-source LLMs on the GSM8k and MATH benchmarks, as well as surpassing several closed-source LLMs including ChatGPT-3.5 and Minerva. MAmmoTH (Yue et al., 2023a) is a series of open-source LLMS specifically developed for general math problem-solving, achieving superior performance on nine mathematical reasoning datasets. Utilizing a novel instruction tuning dataset called MathInstruct, which combines chain-of-thought and program-of-thought rationales, MAmmoTH models demonstrate substantial improvements over existing models. TORA (Gou et al., 2024), a series of Tool-integrated Reasoning Agents, significantly advances mathematical problem-solving by combining natural language reasoning with the use of external computational tools. It markedly outperforms existing open-source models on 10 mathematical reasoning datasets, showcasing notable improvements over both rationale-based and program-based approaches, and introduces innovative training techniques such as output space shaping to enhance model reasoning capabilities. G-LLaVA (Gao et al., 2023c) introduces a significant advancement in geometric problem-solving for LLMs by leveraging a multimodal approach that combines text and image data. This model, utilizing the Geo170K dataset comprising over 170,000 geometric image-caption and question-answer pairs, demonstrates remarkable improvements over GPT-4V on the MathVista benchmark.
@@ 7 Conclusion and Discussion
Legal Considerations for Using LLM Outputs: Importantly, it’s crucial to note the legal implications of utilizing LLM outputs, such as those from ChatGPT [^4] https://openai.com/policies/business-terms, Llama [^5] https://llama.meta.com/llama-downloads/, etc. We strongly advocate compliance with the terms of use specified by the model providers, such as the restrictions on developing competitive products, and so on.